OpenStudy (anonymous):

last one for the night (I hope)... solve for X again: log(7x-2)-log(x-2)=1

OpenStudy (anonymous):

\[\log((7x-2)/(x-2))=1\]

OpenStudy (anonymous):

raise both sides with a base of 10

OpenStudy (anonymous):

\[(7x-2)/(x-2)=10\]

OpenStudy (anonymous):

\[(x-2)*10=7x-2\]

OpenStudy (anonymous):

x= 3 1/3

OpenStudy (anonymous):

Hmm, x should equal 6 I think

OpenStudy (zarkon):

it is 6

OpenStudy (anonymous):

10x - 20 = 7x - 2 3x = 18 x = 6

OpenStudy (anonymous):

dr's process was right, the last calculation was the only thing off

OpenStudy (anonymous):

excellent, thanks to all!

OpenStudy (anonymous):

shoot yes 20 -2 = 18 not 22 LOL

OpenStudy (anonymous):

giving 6